A 12 - diametrical pitch spur gear having 21 teeth rotates at 2000 rpm and drives a pinion at 750 rpm.

Calculate:

a) The number of teeth on the pinion

b) The pitch diameter of the gear

c) The pitch diameter of the pinion


NB: Assume the pitch diameter is measured in inches


1
Expert's answer
2021-02-10T01:12:09-0500

a) 12 = "\frac{15}{Dg}"

"Dg = 1.25" unit

b) "D =\frac{Tc \times Pc}{\pi}"

= "\frac{21 \times 0.262}{\pi}"

= 1.752 unit

c) Tp = "\frac{2000\times 21}{750}"

=56 teeth